# Uncertainty Distillation: Teaching Language Models to Express Semantic Confidence

### v1

@@ Abstract
As large language models (LLMs) are increasingly used for factual question-answering, it becomes more important for LLMs to have the capability to communicate the likelihood that their answer is correct. For these verbalized expressions of uncertainty to be meaningful, they should reflect the error rates at the expressed level of confidence. However, when prompted to express confidence, the error rates of current LLMs are inconsistent with their communicated confidences, highlighting the need for uncertainty quantification methods. Many prior methods calculate lexical uncertainty, estimating a model’s confidence in the specific string it generated. In some cases, however, it may be more useful to estimate semantic uncertainty, or the model’s confidence in the answer regardless of how it is verbalized. We propose a simple procedure, uncertainty distillation, to teach an LLM to verbalize calibrated semantic confidences. Using held-out data to map initial uncertainty estimates to meaningful probabilities, we create examples annotated with verbalized probabilities for supervised fine-tuning. We demonstrate our method yields verbalized confidences that correlate with observed error rates with a small fine-tuned language model as well as with larger instruction-tuned models, and find that our semantic uncertainty correlates well with lexical uncertainty on short answers.
@@
Sophia Hager, David Mueller, Kevin Duh, Nicholas Andrews Johns Hopkins University shager2@jh.edu {dam,noa}@jhu.edu kevinduh@cs.jhu.edu
@@ 1 Introduction
Advances in LLM research have led to instruction-tuned generative models with impressive capabilities on many challenging tasks (OpenAI et al., 2024; Jiang et al., 2023; Dubey et al., 2024). While the flexibility and quality of these models is appealing, they may still hallucinate or give incorrect answers (Rawte et al., 2023; Bai et al., 2024). However, language models do not readily provide an interpretable measure of a model’s likelihood of correctness. LLMs tend to produce poorly-calibrated confidences when prompted to do so, with a propensity for being confidently incorrect Xiong et al. (2024). Furthermore, the elicited confidences may be impacted in unexpected ways by the choice of prompt (Sclar et al., 2023), such as the interpretation of “very confident” being dependent on the wording of the prompt.
There are several other approaches that can be used in lieu of prompting. Models’ token-level probabilities can be used to provide information as a measure of lexical uncertainty, which gives information about the likelihood of a generated string. This is often useful; however, the same fact can be expressed in any number of ways—“Berlin’s the capital of Germany” or “The capital of Germany is Berlin!” or “Die Hauptstadt Deutschlands ist Berlin”—all capturing the same meaning Kuhn et al. (2023). Semantic uncertainty is therefore challenging to capture, as token-level probabilities are influenced by the phrasing of an answer just as much as the semantics of the answer itself. This issue is exacerbated in models employing large vocabularies such as multilingual language models, in language models employing byte or character-level tokenization, or when using LLMs that are prone to producing extraneous outputs Xue et al. (2021); Wang et al. (2024).
@@
We present uncertainty distillation, a scheme for fine-tuning a language model to verbalize uncertainty based on its own internal state. Notably, this is an estimate of semantic rather than lexical uncertainty, as the confidence is estimated from its internal representation without using a separate uncertainty quantification model or information from token-level probabilities. At inference time, models trained using uncertainty distillation efficiently generate a well-calibrated and interpretable statement of confidence in their answers, such as “Berlin is the capital of Germany [high confidence].” [^1] The uncertainty could be expressed in a variety of ways, including using special characters or numeric values. In principle, this approach has several benefits: for instance, semantically equivalent but lexically different predictions can be assigned the same confidence, and a single generation with multiple associated claims can each be assigned different confidences.
Rather than specializing to a specific model or task, uncertainty distillation automatically self-annotates any desired QA dataset with the chosen model’s confidence scores, which may then be used to fine-tune that model to produce verbalized confidences. At a high level, our approach consists of three steps: (1) obtaining semantic uncertainty estimates from the model on held-out calibration data; (2) post-hoc calibrating these into meaningful probabilities; and (3) teaching the model via continued training to output verbalized confidences along with its predictions.
@@
We propose uncertainty distillation, a simple and effective scheme which uses supervised fine-tuning to teach LLMs to output calibrated semantic confidence statements along with predictions. We demonstrate that uncertainty distillation achieves well-calibrated results, both with a single-task fine-tuned model and with instruction-tuned models without initial task-specific training. We compare uncertainty distillation to a powerful lexical uncertainty quantification baseline on simple tasks, and find that we achieve similarly positive results despite operating without lexical uncertainty data at inference time. Finally, we analyze the impact of choices made at different steps of uncertainty distillation, and discuss our recommendations for using uncertainty distillation in practical applications.
@@ 2 Related Work
Generally, calibration refers to the concept that predicted probabilities should align with the probability of correctness (Guo et al., 2017). Mielke et al. (2022) additionally propose the conception of “linguistic calibration”—that models demonstrate uncertainty or doubt through natural language when they are incorrect, determining this uncertainty by using a predictor to determine the likelihood that an answer is correct and considering that to be the model’s uncertainty. There are significant advantages to verbalizing uncertainty: for one, there is relatively low computational overhead to generate several extra tokens, while using a separate calibration model to estimate confidence and then communicate this information to the user requires more computation at inference time Yang et al. (2024). Verbalized confidences are also easily interpretable to an average user regardless of experience or background.
@@
Lexical uncertainty quantification metrics using information from token-level probabilities are commonly used and frequently effective Hu et al. (2023); Malinin and Gales (2021). These probabilities are easily obtainable, do not require additional inference-time compute to generate, and often provide sufficient information for downstream use cases: e.g. error correction in chain of thought Yin et al. (2024), hallucination detection Arteaga et al. (2024), or out-of-distribution data detection Hendrycks et al. (2020). However, there are several disadvantages to lexical uncertainty quantification: it relies on model probabilities which may not be well-calibrated Guo et al. (2017), and is often ineffective on calculating uncertainty of long generations Zhang et al. (2024). The latter in particular may present problems for end users, as models trained using Reinforcement Learning from Human Feedback (RLHF) may be incentivized to produce long outputs Singhal et al. (2024). It is therefore important to consider uncertainty quantification methods that do not rely on token-level probabilities to estimate uncertainty.
@@
Running multiple steps at inference time may provide a better estimate of semantic probability. Xiong et al. (2024) investigate several inference-time strategies which use multiple steps to estimate model uncertainty, such as sampling several answers on the same question or noting if a model changes its answer when prompted with a misleading alternative. While these methods do lead to improvements in LLM calibration, no single intervention consistently emerges as the most successful, and the authors note there is significant scope for improvement. Kuhn et al. (2023) and Farquhar et al. (2024) more explicitly relate this to semantic uncertainty, and find that sampling $m$ predictions from the model and clustering by semantic equivalence results in a robust measure of semantic uncertainty that compares favorably to lexical uncertainty. A major disadvantage of these sampling-based approaches is their increased computational complexity at inference time, however; for instance, Kuhn et al. (2023) recommend sampling at least 10 outputs per question [^2] We compare this approach to our own in Appendix A and find that the additional computation does not consistently outperform uncertainty distillation..
@@ 3.1 Monte Carlo sampling with semantic normalization
Assuming input $x$ and output $y$, we are looking to find $P(y\in Y_{equivalent}\mid x)$, the model’s likelihood of producing this answer or one that is semantically equivalent; however, this would require marginalization over an infinite set of strings $Y$. To make this a tractable problem, we use a Monte Carlo approximation, where our estimate of the models’ predictive distribution improves with $N$, at the expense of additional computational cost. Note however that we do not assume this quantity is a meaningful probability out-of-the-box due to potential overfitting or underfitting of the base model. To diagnose potential miscalibration of the base model as well as correct for it, we may use held-out calibration data that was not seen during training. [^3] We examine the efficacy of our method when this assumption does not hold true in §missing 6.1.
In more detail, to fit a post-hoc calibrator, we need a supervised dataset of datapoints not seen at training time $\{X^{\text{cal}},Y^{\text{cal}}\}$. For each example $x\in X^{\text{cal}}$ we sample $N$ candidate answers $\{\hat{y}_{i}\}_{i=1}^{N}\sim P_{\theta}(Y\mid X=x)$ from a model’s predictive distribution [^4] This model may have been fine-tuned on the specific task as in §missing 4.1 or instruction-tuned as in §missing 4.2.. Before calculating the relative frequency of strings, we apply a normalization function (or set of normalization functions) to consolidate semantically similar outputs. In the short-form QA tasks we consider in §missing 4, we use the simple normalization function of removing punctuation and standardizing capitalization, but for more complex questions other approaches may be required (Huang et al., 2024). For example, Kuhn et al. (2023) use natural language inference to cluster semantically equivalent answers and Tian et al. (2023) use an LLM as a judge of correctness. After consolidating similar events, the relative frequency $f$ of these events is a measure of the LLM’s uncertainty in those events, although this may not be a well-calibrated probability.
@@ 3.2 Post-hoc calibration
In general, neural networks are prone to miscalibration, even when trained using proper losses such as the cross-entropy or log-loss. A common remedy is to apply post-hoc calibration methods, which usually involve some form of regression on the predicted scores to transform them into meaningful probabilities. Two common options are isotonic regression and Platt scaling (sometimes called temperature scaling) (Guo et al., 2017). Our approach uses a model’s predictions on $\{X^{\text{cal}},Y^{\text{cal}}\}$ to diagnose and mitigate badly-calibrated initial model probabilities. We fit an isotonic regression model [^5] Implemented using scikit-learn 1.5.2 on our calibration set by comparing the predicted scores to observed labels. [^6] We discuss the effect this has in §missing 6.3. We compare each prediction $\hat{y}$ with score $f$ to observed events $y$. This yields a calibration map $c:\mathbb{R}\rightarrow[0,1]$ we apply to the relative frequencies of events from samples in the previous step to yield probabilities.
@@ 3.3 Self-annotation and fine-tuning
While correct answers should always be added as training data, appending the confidence scores to incorrect answers may improve the model’s ability to correctly verbalize its own confidence. However, it may also decrease the accuracy of the QA model. We introduce a hyperparameter to control the number of incorrect answers added to the training data. In §missing 6.4, we further discuss the impact of this hyperparameter.
Starting from the sampled model, we perform supervised fine-tuning on these self-annotated targets with verbalized confidences to estimate a second model capable of verbalizing its confidence. If training an instruction-tuned model, we append “Additionally state how confident you are in your answer.” to the preexisting instruction. At inference time, we obtain predictions and verbalized confidences from this new model on held-out test data. We remark that our model incurs little additional cost at inference time, as opposed to other confidence elicitation methods which require inference-time sampling Farquhar et al. (2024); Xiong et al. (2024).
@@ 4 Experimental Setup
We examine the efficacy of uncertainty distillation in two settings, with consistent metrics and baseline. In the first, we consider uncertainty quantification on a model that has been fine-tuned to perform a single task. In the second, we consider a multi-task instruction-tuned model on a set of unseen QA tasks, and examine whether our method can achieve reasonable estimates of uncertainty without the model first being fine-tuned on those tasks.
4.1 Supervised fine-tuning experiment

### v2

# Can Language Models Verbalize Well-Defined Notions of Uncertainty?

@@ Abstract
As large language models (LLMs) are increasingly used for factual question-answering, it becomes more important for LLMs to have the capability to communicate the likelihood that their answer is correct. For these verbalized expressions of uncertainty to be meaningful, they should reflect the error rates at the expressed level of confidence. However, when prompted to express confidence, the error rates of current LLMs are inconsistent with their communicated confidences, highlighting the need for uncertainty quantification methods. Many prior methods calculate lexical uncertainty, estimating a model’s confidence in the specific string it generated. In some cases, however, it may be more useful to estimate semantic uncertainty, or the model’s confidence in the answer regardless of how it is verbalized. We propose a simple procedure, uncertainty distillation, to teach an LLM to verbalize calibrated semantic confidences. Using held-out data to map initial uncertainty estimates to meaningful probabilities, we create examples annotated with verbalized probabilities for supervised fine-tuning. We compare uncertainty distillation to several strong baselines, and find that our method yields verbalized confidences that correlate well with observed error rates.
@@
Sophia Hager, David Mueller, Kevin Duh, Nicholas Andrews Johns Hopkins University {shager2,noa}@jhu.edu
@@ 1 Introduction
Advances in LLM research have led to instruction-tuned generative models with impressive capabilities on many challenging tasks (OpenAI et al., 2024; Jiang et al., 2023; Dubey et al., 2024). While the flexibility and quality of these models is appealing, they may still hallucinate or give incorrect answers (Rawte et al., 2023; Bai et al., 2024). However, language models do not readily provide an interpretable measure of a model’s likelihood of correctness. LLMs tend to produce poorly-calibrated confidences when prompted to do so, and are often confidently incorrect Xiong et al. (2024). Furthermore, the elicited confidences may be impacted in unexpected ways by the choice of prompt (Sclar et al., 2023), such as the interpretation of “very confident” being dependent on the wording of the prompt.
There are several other approaches as an alternative to prompting. Models’ token-level probabilities can be used to provide information as a measure of lexical uncertainty, which gives information about the likelihood of a generated string. This is often useful; however, the same fact can be expressed in any number of ways—“Berlin’s the capital of Germany” or “The capital of Germany is Berlin!” or “Die Hauptstadt Deutschlands ist Berlin”—all capturing the same meaning Kuhn et al. (2023). Semantic uncertainty is therefore challenging to capture, as token-level probabilities are influenced by the phrasing of an answer just as much as the semantics of the answer itself. This issue is particularly challenging for models employing large vocabularies such as multilingual language models, language models employing byte or character-level tokenization, or when using LLMs that are prone to producing extraneous outputs Xue et al. (2021); Wang et al. (2024).
@@
We present uncertainty distillation, a scheme for fine-tuning a language model to verbalize uncertainty based on its own internal state. Notably, uncertainty distillation teaches models to estimate their semantic—rather than lexical—uncertainty, as the distilled confidences are estimated from the probabilities of semantically normalized outputs, rather than relying on token-level probabilities. At inference time, models trained using uncertainty distillation efficiently generate a well-calibrated and interpretable statement of confidence in their answers, such as “Berlin is the capital of Germany [high confidence].” [^1] The uncertainty could be expressed in a variety of ways, including using special characters or numeric values. Our approach enables semantically equivalent but lexically different predictions to be assigned the same confidence, and a single generation with multiple claims can each be assigned different confidences. Additionally, uncertainty distillation is computationally inexpensive at inference time, generating only a handful of additional tokens. And compared to methods such as P(IK) Farquhar et al. (2024), we do not require a separate uncertainty network; our approach uses standard supervised fine-tuning recipes for LLMs.
Rather than specializing to a specific model or task, uncertainty distillation involves self-annotation of any desired QA dataset with the base model’s calibrated uncertainties, which are then used to fine-tune that model to produce verbalized confidences. At a high level (Figure 1), our approach consists of three steps: (1) obtaining semantic uncertainty estimates from the model on held-out calibration data; (2) post-hoc calibrating these into meaningful probabilities; and (3) teaching the model via supervised fine-tuning to output verbalized confidences along with its predictions.
@@
  • We propose uncertainty distillation, a simple yet effective scheme which uses supervised fine-tuning to teach LLMs to output calibrated semantic confidence statements along with their predictions.
  • We demonstrate that uncertainty distillation achieves easily interpretable results and compares favorably to several powerful baselines.
  • We analyze the impact of choices made when implementing uncertainty distillation.
@@ 2 Related Work
Generally, calibration refers to the concept that predicted probabilities should align with the probability of correctness (Guo et al., 2017). Mielke et al. (2022) additionally propose the conception of “linguistic calibration"—that models demonstrate uncertainty or doubt through natural language when they are incorrect, determining this uncertainty by using a predictor to determine the likelihood that an answer is correct and considering that to be the model’s uncertainty. There are significant advantages to verbalizing uncertainty: for one, there is relatively low computational overhead to generate several extra tokens, while using a separate calibration model to estimate confidence and then communicate this information to the user requires more computation at inference time Yang et al. (2024). Verbalized confidences are also readily interpretable to an LLM when reasoning about uncertainty, or to an average end-user regardless of experience or background.
@@
Lexical uncertainty quantification metrics using information from token-level probabilities are commonly used and frequently effective Hu et al. (2023); Malinin and Gales (2021). These probabilities are easily obtainable, do not require additional inference-time compute to generate, and often provide sufficient information for downstream use cases: e.g. error correction in chain of thought Yin et al. (2024), hallucination detection Arteaga et al. (2024), or out-of-distribution data detection Hendrycks et al. (2020b). However, there are several disadvantages to lexical uncertainty quantification: it relies on model probabilities which may not be well-calibrated Guo et al. (2017), and is often ineffective on calculating uncertainty of long generations Zhang et al. (2024). The latter, in particular, may present problems for end users, as models trained using Reinforcement Learning from Human Feedback (RLHF) are often incentivized to produce long outputs Singhal et al. (2024). It is therefore important to consider uncertainty quantification methods that do not rely on token-level probabilities to estimate uncertainty.
@@
Running multiple steps at inference time may provide a better estimate of semantic probability. Xiong et al. (2024) investigate several inference-time strategies which use multiple steps to estimate model uncertainty, such as sampling several answers on the same question or noting if a model changes its answer when prompted with a misleading alternative. While these methods do lead to improvements in LLM calibration, no single intervention consistently emerges as the most successful, and the authors note there is significant scope for improvement. Kuhn et al. (2023) and Farquhar et al. (2024) more explicitly relate this to semantic uncertainty, and find that sampling $m$ predictions from the model and clustering by semantic equivalence results in a robust measure of semantic uncertainty that compares favorably to lexical uncertainty. A major disadvantage of these sampling-based approaches is their increased computational complexity at inference time, however; for instance, the semantic clustering approach of Farquhar et al. (2024), which we compare to in our experiments, requires 20 samples at inference time and calls to a separate entailment model.
@@ 3.1 Monte Carlo sampling with semantic normalization
Assuming input $x$ and output $y$, we are looking to find $P(y\in Y_{\text{equivalent}}\mid x)$, the model’s likelihood of producing this answer or one that is semantically equivalent; however, this would require marginalization over an infinite set of strings $Y$. To make this a tractable problem, we use a Monte Carlo approximation, where our estimate of the models’ predictive distribution improves with $N$, at the expense of additional computational cost. Note however that we do not assume this quantity is a meaningful probability out-of-the-box due to potential overfitting or underfitting of the base model. To diagnose potential miscalibration of the base model as well as correct for it, we may use held-out calibration data that was not seen during training. [^2] We examine the efficacy of our method when this assumption does not hold true in §missing 6.1.
In more detail, to fit a post-hoc calibrator, we need a supervised dataset of datapoints not seen at training time $\{X^{\text{cal}},Y^{\text{cal}}\}$. For each example $x\in X^{\text{cal}}$ we sample $N$ candidate answers $\{\hat{y}_{i}\}_{i=1}^{N}\sim P_{\theta}(Y\mid X=x)$ from a model’s predictive distribution [^3] This model may have been fine-tuned on the specific task as in Appendix A or instruction-tuned as in §missing 4.2.. Before calculating the relative frequency of strings, we apply a normalization function (or set of normalization functions) to consolidate semantically similar outputs. In the short-form QA tasks we consider in §missing 4, we use the simple normalization function of isolating a multiple choice answer using tags, removing punctuation and standardizing capitalization, but for more complex questions other approaches may be required (Huang et al., 2024). For example, Kuhn et al. (2023) use natural language inference to cluster semantically equivalent answers and Tian et al. (2023) use an LLM as a judge of correctness. After consolidating similar events, the relative frequency $f$ of these events is a measure of the LLM’s uncertainty in those events, although this may not be a well-calibrated probability.
@@ 3.2 Post-hoc calibration
In general, neural networks are prone to miscalibration, even when trained using proper losses such as cross-entropy. A common remedy is to apply post-hoc calibration methods, which usually involve some form of regression on the predicted scores to transform them into meaningful probabilities. Specifically, we post-hoc calibrate the relative frequencies of each semantic cluster found in the previous step. Two common options for post-hoc calibration are isotonic regression and Platt scaling (sometimes called temperature scaling) (Guo et al., 2017). Our approach uses a model’s predictions on $\{X^{\text{cal}},Y^{\text{cal}}\}$ to diagnose and mitigate badly-calibrated initial model probabilities. We fit an isotonic regression model [^4] Implemented using scikit-learn 1.5.2 on our calibration set by comparing the predicted scores to observed labels. [^5] We discuss post-hoc calibration further in Appendix D. We compare each prediction $\hat{y}$ with score $f$ to observed events $y$. This yields a calibration map $c:\mathbb{R}\rightarrow[0,1]$ we apply to the relative frequencies of events from samples in the previous step to yield probabilities.
@@ 3.3 Self-annotation and fine-tuning
While correct answers should always be added as training data, appending the confidence scores to incorrect answers may improve the model’s ability to correctly verbalize its own confidence. However, it may also decrease the accuracy of the QA model. We introduce a hyperparameter to control the number of incorrect answers added to the training data. In §missing 6.2, we further investigate the impact of this hyperparameter.
Starting from the sampled model, we perform supervised fine-tuning on these self-annotated targets with verbalized confidences to estimate a second model capable of verbalizing its confidence. If training an instruction-tuned model, we append an additional instruction such as "Additionally state how confident you are in your answer." to the preexisting instruction [^6] See Appendix C for details on the specific prompts used in each experiment.. If a reasoning trace has been generated during sampling, we randomly select a reasoning trace to add to the target answer from all possible options. At inference time, we obtain predictions and verbalized confidences from this new model on held-out test data. We remark that our model incurs little additional cost at inference time, as opposed to other confidence elicitation methods which require inference-time sampling Farquhar et al. (2024); Xiong et al. (2024).
@@ 4 Experimental Setup
Dataset Model Method AUROC Accuracy High Accuracy
MMLU Ministral-8B Uncertainty distillation 0.693 0.601 0.766
Lexical baseline 0.627 0.551 0.555
Prompting 0.587 0.637 0.643
P(IK) 0.555 0.518 0.553
P(True) 0.471 0.585 0.583
Semantic clustering 0.667 0.577 0.821
Llama-3B Uncertainty distillation 0.743 0.532 0.759
Lexical baseline 0.644 0.511 0.600
Prompting 0.548 0.613 0.647
P(IK) 0.669 0.515 0.681
P(True) 0.550 0.554 0.558
Semantic clustering 0.646 0.560 0.727
SocialIQA Ministral-8B Uncertainty distillation 0.632 0.712 0.800
Lexical baseline 0.600 0.738 0.760
Prompting 0.539 0.721 0.738
P(IK) 0.587 0.614 0.656
P(True) 0.491 0.712 0.710
Semantic clustering 0.603 0.659 0.780
Llama-3B Uncertainty distillation 0.784 0.653 0.833
Lexical baseline 0.531 0.673 0.687
Prompting 0.545 0.685 0.712
P(IK) 0.595 0.642 0.811
P(True) 0.505 0.681 0.682
Semantic clustering 0.601 0.675 0.758
Table 1: AUROC and accuracy metrics for our large models and datasets. We find that uncertainty distillation leads to increased AUROC and accuracy in high-confidence categories, albeit with a small decrease in overall accuracy. Accuracy is the overall accuracy, and High Accuracy is the accuracy for the most confident predictions. We find that uncertainty distillation with one generation achieves High Accuracy better than or similar to semantic clustering, which requires 20 samples at inference time.
We examine the efficacy of uncertainty distillation in two settings. First, we demonstrate the success of uncertainty quantification with large language models on several standard QA benchmarks. In the second, we consider two smaller multi-task instruction-tuned models on a set of unseen QA tasks and analyze the impact of various choices made during uncertainty distillation. In particular, this setting allows us to examine the effect of using models which have been pretrained on our calibration set for uncertainty distillation.
4.1 Uncertainty Distillation on LLMs
Datasets
We demonstrate uncertainty distillation using two multiple-choice question answering datasets, the Massive Multitask Language Understanding benchmark (MMLU) Hendrycks et al. (2020a) and the Social Interaction Question Answering dataset (SocialIQA) Sap et al. (2019). MMLU is a multitask dataset consisting of multiple choice questions over 57 subjects such as high school psychology or formal logic. We take a subset of 20,000 questions from the training set to act as our calibration data, a subset of 500 questions from the validation set to act as our validation data, and a subset of 2,000 quesions from the test set to act as our test data. SocialIQA is a dataset consisting of question/answer pairs about social situations. We take a subset of 20,000 questions from the training set to act as our calibration data, a subset of 500 questions from the training set to act as our validation data, and use the existing validation split as our test data. For both datasets we set $N=100$, i.e. we take 100 samples per question to construct our initial Monte Carlo estimate of confidence.
Models
We validate uncertainty distillation on these datasets using two modern instruction-tuned LLMs, Llama-3.2-3B-Instruct Dubey et al. (2024) and Ministral-8B-Instruct-2410 Jiang et al. (2023). When performing uncertainty distillation with Ministral-8B, we use LoRA Hu et al. (2021).
Baselines
For the Lexical baseline, we extract token-level probabilities from the language model on our training/calibration split [^7] As we do not have an initial fine-tuning step, these are equivalent. and use this to train an isotonic regression model to calibrate the average token-level probability for each answer. [^8] We use the average probability rather than the sequence probability to normalize over different lengths, as Kuhn et al. (2023) find this improves performance. Prompting: To measure the model’s ability to verbalize its confidence prior to uncertainty distillation we prompt the model to output its own confidence in its answer without any additional training. We report this baseline for these models, and discuss the prompts used in Appendix C. We also compare to P(IK) from Farquhar et al. (2024) which learns a mapping from hidden states to uncertainty scores, and P(True) from Kadavath et al. (2022). Finally, we compare to the Semantic Clustering approach from Farquhar et al. (2024). Both P(True) and Semantic Clustering generate 20 samples from the model to compute uncertainty scores, unlike our approach which uses a single generation.
4.2 Targeted analysis
In §missing 3, we assume that we have access to held-out calibration data. However, due to the opacity and size of pretraining datasets, it is increasingly challenging to guarantee that this assumption holds. We therefore test uncertainty distillation in a setting where we can know with certainty whether the calibration set is in a model’s pretraining data. We examine this question and the impact of adding varying numbers of incorrect answers during uncertainty distillation in §missing 6.
Dataset
In this setting, we use the Super-NaturalInstructions dataset (SNI; Wang et al., 2022). We select $15$ English Q&A tasks with short-form answers. We focus on Q&A tasks for which a single correct answer exists (e.g. multiple choice problems, short-form span extraction, math problems, etc.) and thus for which correctness of a model’s prediction can reliably and efficiently be computed after normalizing lexical forms without resorting to methods such as LLM verification. We use 1,000 samples to obtain our Monte Carlo estimate of confidence (see Appendix B for details on how number of samples affects successful confidence estimation).
Models
We perform uncertainty distillation on FLAN-T5 Chung et al. (2022), an instruction-tuned model trained on a dataset containing the SNI tasks. Importantly, we not only verify that Flan-T5 has been instruction-tuned on our tasks, but has seen samples from the calibration set of our test tasks. This allows us to investigate the effect of data contamination on calibration of verbalized confidences.
To construct a similar model which has not seen our calibration data, we instruction-tune a T5-Large model on a remaining subset of the English tasks in the SNI dataset, making sure to explicitly hold out the $15$ tasks we use in our uncertainty distillation experiments. The result is an instruction-tuned model which we refer to as Instruct-T5, capable of performing our target Q&A tasks without having seen these tasks during training. In other words, the samples we obtain from this model do not require Instruct-T5 to be pre-trained on that specific task. See Appendix E for more details on our data selection and instruction-tuning. We train and evaluate uncertainty distillation on the combined dataset of these tasks and report the performance over the metrics described in §missing 4.3.
Baselines
We report a comparison to the lexical baseline described above in order to provide context for the performance of the small models.
4.3 Metrics
We report the area under the receiver operating characteristic curve (AUROC), [^9] Calculated using scikit-learn 1.5.2 which represents the probability that a randomly chosen correct answer will be in a higher-confidence bin than a randomly chosen incorrect answer. This metric is well established in previous literature (see e.g., Hu et al. (2023)), and compares the relative rather than absolute probabilities, which allows us to use it effectively with discrete verbalized confidences. [^10] We do not report Expected Calibration Error (ECE) (Nixon et al., 2019), as it requires comparing a continuous score to the prediction’s true label, while we have five discrete bins without an obvious corresponding numerical score. Furthermore, it requires the choice of several hyperparameters which can have a large impact on performance. Baseline methods that return a continuous score are binned to five categories to represent converting to a comparable verbalized confidence. For all methods, we plot the percentage of accurate answers in each bin to examine if confidence corresponds well with accuracy. We also report overall model accuracy, to evaluate the tradeoff between accuracy and calibration. Finally, we report high accuracy, accuracy of high-confidence predictions (in “very high” and “high” bins). As an established use-case for verbalized confidences is to reject lower-confidence predictions, this provides information about how useful the LLM’s predictions are. [^11] The fact that high accuracy is not perfect also highlights a risk of confidence estimation: namely, that it increases trust in an answer that still may be incorrect.
Figure 2: Average accuracy within each confidence bin for our experiments with Llama (Mistral results in Figure 5). We find that our confidence bins correspond well with accuracy within the bin, while our baselines may not exhibit similar correspondence. We do not plot bins with fewer than 10 samples.
5 Results and discussion
Figure 2 shows some of our results comparing uncertainty distillation to the lexical uncertainty baseline in terms of average accuracies in each confidence bin [^12] We present the remaining two settings in Appendix F.. In plots like this, the goal is to observe a diagonal trend line where outputs reported to have high confidence indeed have high accuracy, and those in the low confidence bins have lower accuracy. We find that the verbalized confidences produced by uncertainty distillation are highly interpretable, with high correspondence between accuracy of answers within a bin and that bins confidence. In contrast, confidence scores generated by the baselines may not correspond well with the actual accuracies within that bin. For instance, accuracy within the lowest confidence bin for the prompting baseline is 0.684 with Llama-3B on SocialIQA, while accuracy within the highest confidence bin is 0.651.
Table 1 summarizes these plots in terms of AUROC score. AUROC is consistently higher with uncertainty distillation than with other methods. We conclude that uncertainty distillation is an effective method for estimating confidence in an answer. AUROC is consistently highest for uncertainty distillation with both models on both datasets. In particular, we note that uncertainty distillation consistently achieves higher AUROC than semantic clusteringKuhn et al. (2023), despite semantic clustering requiring 20 samples and a computationally intensive clustering step at inference time: for instance, uncertainty distillation achieves AUROC of 0.784 with Llama-3B on SocialIQA, while semantic clustering achieves AUROC of 0.601.
The table also reports the accuracy of the highest confidence bin and the overall accuracy across all bins. While AUROC is the main metric for assessing calibration performance, accuracy is also useful for understanding the nuances of the result. We find that uncertainty distillation does not lead to notable drops in overall accuracy, and that accuracy in the highest bins increases dramatically. Uncertainty distillation achieves the best high accuracy in all but one case, Ministral-8B on MMLU. In this case, semantic clustering achieves high accuracy of 0.821 compared to uncertainty distillation’s high accuracy of 0.766. However, we again note the comparative computational complexity of semantic clustering, as well as the fact that uncertainty distillation’s high accuracy is significantly higher than the next highest, prompting with 0.643.
6 Analysis
Model Method AUROC Overall Accuracy High Accuracy
Instruct-T5 Uncertainty distillation 0.751 0.449 0.839
Lexical baseline 0.667 0.387 0.754
FLAN-T5 Uncertainty distillation 0.873 0.614 0.875
Lexical baseline 0.892 0.657 0.912
Table 2: AUROC and accuracy metrics when using FLAN-T5, which does not have an unseen calibration set. We find that while uncertainty distillation outperforms our lexical baseline with a model with an unseen calibration set, it does not outperform the baseline on FLAN-T5, which was instruction-tuned on the data previously.
6.1 Assumption of calibration set
We compare the performance of FLAN-T5, which has been instruction-tuned on the calibration set, with the performance of Instruct-T5, which has not, in Table 2. We find that while uncertainty distillation still produces meaningful confidence bins for FLAN-T5, it no longer outperforms lexical uncertainty. We conclude that uncertainty distillation works in the absence of held-out calibration data, but not as effectively as token-level probabilities, which are likely well-calibrated due to the model’s previous training on these examples. We discuss results for these two models further in §missing 6.2 and Appendix D, and find that the behavior of FLAN-T5 differs significantly from results on models where we have an unseen calibration set.
6.2 Adding incorrect examples
While adding incorrect examples into the training data has the potential to provide more examples at different levels of confidences, it also is likely to increase the likelihood that a model generates an incorrect answer. To demonstrate this effect, in Table 3, we show the AUROC and accuracy for models trained with different amounts of incorrect samples. With Instruct-T5, we find that adding only two incorrect samples per correct sample dramatically increases AUROC while decreasing accuracy. While this would seem to indicate a fundamental tradeoff between accuracy and calibration, we find that the same is not as obviously true for FLAN-T5; while the accuracy may decrease and AUROC may increase, the effects are not as significant as they are for Instruct-T5. One possible interpretation of this is that its predictions are shaped by the fact that the data was included in its instruction-tuning corpus, leading to less dramatic shifts when trained.
While adding incorrect samples may improve AUROC, it increases the number of training examples by a factor of the number of incorrect examples added (e.g. a training set with 100 examples would train on 100 augmented answers with 0 incorrect examples added, 200 augmented answers with one incorrect example added, etc.) This leads to increased compute at training time. For this reason, in addition to the decreased accuracy, we recommend adding a low number of incorrect examples to the training dataset, and in our main experiments limit to at most one incorrect answer per question.
0 1 2 3
Instruct-T5
AUROC 0.723 0.737 0.751 0.757
Accuracy 0.529 0.486 0.449 0.447
FLAN-T5
AUROC 0.868 0.876 0.873 0.883
Accuracy 0.609 0.620 0.614 0.611
Table 3: AUROC of models trained with varying numbers of incorrect examples allowed per question. There is a general trend towards increasing AUROC and decreasing accuracy when incorrect examples are included, although this is less pronounced for FLAN-T5.
7 Conclusion
Findings
We find that uncertainty distillation leads to improved estimates of uncertainty in comparison to many strong baselines, including baselines that require considerably more samples at inference-time. One high-level takeaway is that with small models there appears to be a tradeoff between an LLM’s ability to predict its own confidence and overall model accuracy, but that this effect is less obvious with increasing model sizes. In our small-scale analysis, interventions that improve AUROC decrease accuracy and vice versa; however, with larger models we do not note as noticeable a decrease in accuracy compared to our baselines. Overall, we view our contribution as a significant step towards LLMs that can reliably reason about uncertainty, without requiring any auxiliary models or incurring additional inference-time compute.
Future work
While we focus on QA tasks, our method could be applied to tasks outside simple QA through the use of LLM verifiers to calculate binary correctness. Future work may also investigate the robustness of the model’s internal representation of uncertainty to domain shifts—do the confidence scores remain interpretable and well-calibrated on similar unseen tasks without further training? Looking beyond these immediate questions, LLMs that are able to verbalize meaningful confidences, for example thanks to our method, may be useful in a variety of applications requiring reasoning about uncertainty, such as medical diagnosis.
Limitations
Our experiments focus on established QA tasks which admit straightforward ways to assess correctness. In principle, our approach generalizes to more complex tasks involving longer-form generations, for example using an LLM verifier to establish correctness; we leave it as future work to experiment in these settings. Separately, the proposed approach may be useful in cases where a single generation involves multiple distinct claims that each need to be associated with distinct confidences. Future work should identify appropriate datasets to evaluate multi-claim uncertainty estimation. Finally, our experiments do not include models larger than 8 billion parameters due to compute limitations; however, we hope that our findings will encourage further study into uncertainty distillation for larger LLMs and in more general settings.
Appendix A Uncertainty distillation on supervised fine-tuned models
We here examine uncertainty distillation’s efficacy when performed on a small fine-tuned model, rather than large instruction-tuned models.